find distance between (-3,-2) and (1,7)
Answer by checkley71(8403)

ON GRAPH PAPER LOCATE THESE 2 POINTS.
THEN DRAW A RIGHT TRIANGLE USING THE X & Y AXIS.
THE DISTANCE BETWEEN THESE 2 POINTS IS THE HYPOTENUSE OF THIS TRIANGLE 7 THE FORMULA IS A^2+B^2=C^2
NOW WE FIND THE LENGTH OF THE X & Y AXIS SIDES.
X=1+3=4
Y=7=2=9
4^2+9^2=C^2
16+81=C^2
C^2=97
C=SQRT97
C=9.85 IS THE HYPOTENUSE & ALSO THE DISTANCE BETWEEN THESE 2 POINTS.